A man can hit a target once in 4 shots. If he fires 4 shots in succession, what is the probability that he will hit his target?
সমাধান
সঠিক উত্তর: Option A
Probability of a hit in one shot = 1/4
Probability of a miss in one shot = 1-1/4=3/4
1 - P(misses 1st time AND misses 2nd time AND misses 3rd time AND misses 4th time)
"AND" means "MULTIPLY":
1 - P(misses 1st time)*P(misses 2nd time)*P(misses 3rd time)*P(misses 4th time) so probability=1-(3/4X3/4X3/4x3/4)=1-81/256=175/256